# Agentic Retrieval-Augmented Generation: A Survey on Agentic RAG

### 15 Jan 2025 v1

@@ Abstract
Agentic Retrieval-Augmented Generation (Agentic RAG) transcends these limitations by embedding autonomous AI agents into the RAG pipeline. These agents leverage agentic design patterns—reflection, planning, tool use, and multi-agent collaboration—to dynamically manage retrieval strategies, iteratively refine contextual understanding, and adapt workflows to meet complex task requirements. This integration enables Agentic RAG systems to deliver unparalleled flexibility, scalability, and context-awareness across diverse applications.
This survey provides a comprehensive exploration of Agentic RAG, beginning with its foundational principles and the evolution of RAG paradigms. It presents a detailed taxonomy of Agentic RAG architectures, highlights key applications in industries such as healthcare, finance, and education, and examines practical implementation strategies. Additionally, it addresses challenges in scaling these systems, ensuring ethical decision-making, and optimizing performance for real-world applications, while providing detailed insights into frameworks and tools for implementing Agentic RAG  [^1] GitHub link: https://github.com/asinghcsu/AgenticRAG-Survey.
@@ 1 Introduction
The evolution of agents [12] has further augmented the capabilities of AI systems. Modern agents, including LLM-powered and mobile agents [13], are intelligent entities capable of perceiving, reasoning, and autonomously executing tasks. These agents leverage Agentic workflow patterns, such as reflection [14], planning [15], tool use, and multi-agent collaboration [16], enabling them to manage dynamic workflows and tackle intricate problems.
The convergence of RAG and agentic intelligence has given rise to Agentic Retrieval-Augmented Generation (Agentic RAG) [17], a paradigm that integrates agents into the RAG pipeline. Agentic RAG enables dynamic retrieval strategies, contextual understanding, and iterative refinement [18], allowing for adaptive and efficient information processing. Unlike traditional RAG, Agentic RAG employs autonomous agents to orchestrate retrieval, filter relevant information, and refine responses, excelling in scenarios requiring precision and adaptability.
@@
The structure of this paper is as follows: Section 2 introduces RAG and its evolution, highlighting the limitations of traditional approaches. Section 3 elaborates on the principles of agentic intelligence and agentic patterns. Section 4 provides a taxonomy of Agentic RAG systems, including single-agent, multi-agent, and graph-based frameworks. Section 5 examines applications of Agentic RAG, while Section 6 discusses implementation tools and frameworks. Section 7 focuses on benchmarks and dataset, and Section 8 concludes with future directions for Agentic RAG systems.
@@ 3.1.2 Planning
Planning [15] is a key design pattern in agentic workflows that enables agents to autonomously decompose complex tasks into smaller, manageable subtasks. This capability is essential for multi-hop reasoning and iterative problem-solving in dynamic and uncertain scenarios as shown in Figure 8.
@@
Figure 8: An Overview of Agentic Planning
@@ 3.1.3 Tool Use
Tool Use enables agents to extend their capabilities by interacting with external tools, APIs, or computational resources as illustrated in 9. This pattern allows agents to gather information, perform computations, and manipulate data beyond their pre-trained knowledge. By dynamically integrating tools into workflows, agents can adapt to complex tasks and provide more accurate and contextually relevant outputs.
Figure 9: An Overview of Tool Use
@@ 3.1.4 Multi-Agent
These patterns are foundational to the success of Agentic RAG systems, enabling them to dynamically adapt retrieval and generative workflows to meet the demands of diverse and dynamic environments. By leveraging these patterns, agents can handle iterative, context-aware tasks that exceed the capabilities of traditional RAG systems.
4 Taxonomy of Agentic RAG Systems
@@
4.1 Single-Agent Agentic RAG: Router

### 03 Feb 2025 v2

@@ Abstract
Agentic Retrieval-Augmented Generation (Agentic RAG) transcends these limitations by embedding autonomous AI agents into the RAG pipeline. These agents leverage agentic design patterns reflection, planning, tool use, and multi-agent collaboration to dynamically manage retrieval strategies, iteratively refine contextual understanding, and adapt workflows through clearly defined operational structures ranging from sequential steps to adaptive collaboration. This integration enables Agentic RAG systems to deliver unparalleled flexibility, scalability, and context-awareness across diverse applications.
This survey provides a comprehensive exploration of Agentic RAG, beginning with its foundational principles and the evolution of RAG paradigms. It presents a detailed taxonomy of Agentic RAG architectures, highlights key applications in industries such as healthcare, finance, and education, and examines practical implementation strategies. Additionally, it addresses challenges in scaling these systems, ensuring ethical decision-making, and optimizing performance for real-world applications, while providing detailed insights into frameworks and tools for implementing Agentic RAG  [^1] GitHub link: https://github.com/asinghcsu/AgenticRAG-Survey. The GitHub link for this survey is available at: https://github.com/asinghcsu/AgenticRAG-Survey.
@@ 1 Introduction
The evolution of agents [12] has significantly enhanced the capabilities of AI systems. Modern agents, including LLM-powered and mobile agents [13], are intelligent entities capable of perceiving, reasoning, and autonomously executing tasks. These agents leverage agentic patterns, such as reflection [14], planning [15], tool use, and multi-agent collaboration [16], to enhance decision-making and adaptability.
Furthermore, these agents employ agentic workflow patterns [12, 13], such as prompt chaining, routing, parallelization, orchestrator-worker models, and evaluator-optimizer , to structure and optimize task execution. By integrating these patterns, Agentic RAG systems can efficiently manage dynamic workflows and address complex problem-solving scenarios. The convergence of RAG and agentic intelligence has given rise to Agentic Retrieval-Augmented Generation (Agentic RAG) [14], a paradigm that integrates agents into the RAG pipeline. Agentic RAG enables dynamic retrieval strategies, contextual understanding, and iterative refinement [15], allowing for adaptive and efficient information processing. Unlike traditional RAG, Agentic RAG employs autonomous agents to orchestrate retrieval, filter relevant information, and refine responses, excelling in scenarios requiring precision and adaptability. The overview of Agentic RAG is in figure 1.
@@
The structure of this paper is as follows: Section 2 introduces RAG and its evolution, highlighting the limitations of traditional approaches. Section 3 elaborates on the principles of agentic intelligence and agentic patterns. Section 4 elaborates agentic workflow patterns. Section 5 provides a taxonomy of Agentic RAG systems, including single-agent, multi-agent, and graph-based frameworks. Section 6 examines applications of Agentic RAG, while Section 7 discusses implementation tools and frameworks. Section 8 focuses on benchmarks and dataset, and Section 9 concludes with future directions for Agentic RAG systems.
Figure 1: An Overview of Agentic RAG
@@ 3.0.2 Planning
Planning [24] is a key design pattern in agentic workflows that enables agents to autonomously decompose complex tasks into smaller, manageable subtasks. This capability is essential for multi-hop reasoning and iterative problem-solving in dynamic and uncertain scenarios as shown in Figure 9(a).
@@
(a) An Overview of Agentic Planning
(b) An Overview of Tool Use
Figure 9: Overview of Agentic Planning and Tool Use
@@ 3.0.3 Tool Use
Tool Use enables agents to extend their capabilities by interacting with external tools, APIs, or computational resources as illustrated in 9(b). This pattern allows agents to gather information, perform computations, and manipulate data beyond their pre-trained knowledge. By dynamically integrating tools into workflows, agents can adapt to complex tasks and provide more accurate and contextually relevant outputs.
@@ 3.0.4 Multi-Agent
These design patterns form the foundation for the success of Agentic RAG systems. By structuring workflows—from simple, sequential steps to more adaptive, collaborative processes—these patterns enable systems to dynamically adapt their retrieval and generative strategies to the diverse and ever-changing demands of real-world environments. Leveraging these patterns, agents are capable of handling iterative, context-aware tasks that significantly exceed the capabilities of traditional RAG systems.
4 Agentic Workflow Patterns: Adaptive Strategies for Dynamic Collaboration
Agentic workflow patterns, [12, 13] structure LLM-based applications to optimize performance, accuracy, and efficiency. Different approaches are suitable depending on task complexity and processing requirements.
4.1 Prompt Chaining: Enhancing Accuracy Through Sequential Processing
Prompt chaining [12, 13] decomposes a complex task into multiple steps, where each step builds upon the previous one. This structured approach improves accuracy by simplifying each subtask before moving forward. However, it may increase latency due to sequential processing.
Figure 11: Illustration of Prompt Chaining Workflow
When to Use: This workflow is most effective when a task can be broken down into fixed subtasks, each contributing to the final output. It is particularly useful in scenarios where step-by-step reasoning enhances accuracy.
Example Applications:
  • Generating marketing content in one language and then translating it into another while preserving nuances.
  • Structuring document creation by first generating an outline, verifying its completeness, and then developing the full text.
4.2 Routing:Directing Inputs to Specialized Processes
Routing [12, 13] involves classifying an input and directing it to an appropriate specialized prompt or process. This method ensures distinct queries or tasks are handled separately, improving efficiency and response quality.
Figure 12: Illustration Routing Workflow
When to Use: Ideal for scenarios where different types of input require distinct handling strategies, ensuring optimized performance for each category.
Example Applications:
  • Directing customer service queries into categories such as technical support, refund requests, or general inquiries.
  • Assigning simple queries to smaller models for cost efficiency, while complex requests go to advanced models.
4.3 Parallelization: Speeding Up Processing Through Concurrent Execution
Parallelization [12, 13] divides a task into independent processes that run simultaneously, reducing latency and improving throughput. It can be categorized into sectioning (independent subtasks) and voting (multiple outputs for accuracy).
Figure 13: Illustration of Parallelization Workflow
When to Use: Useful when tasks can be executed independently to enhance speed or when multiple outputs improve confidence.
Example Applications:
  • Sectioning: Splitting tasks like content moderation, where one model screens input while another generates a response.
  • Voting: Using multiple models to cross-check code for vulnerabilities or analyze content moderation decisions.
4.4 Orchestrator-Workers: Dynamic Task Delegation
This workflow [12, 13] features a central orchestrator model that dynamically breaks tasks into subtasks, assigns them to specialized worker models, and compiles the results. Unlike parallelization, it adapts to varying input complexity.
Figure 14: Illustration of Orchestrator-Workers Workflow
When to Use: Best suited for tasks requiring dynamic decomposition and real-time adaptation, where subtasks are not predefined.
Example Applications:
  • Automatically modifying multiple files in a codebase based on the nature of requested changes.
  • Conducting real-time research by gathering and synthesizing relevant information from multiple sources.
4.5 Evaluator-Optimizer: Refining Output Through Iteration
The evaluator-optimizer [12, 13] workflow iteratively improves content by generating an initial output and refining it based on feedback from an evaluation model.
Figure 15: Illustration of Evaluator-Optimizer Workflow
When to Use: Effective when iterative refinement significantly enhances response quality, especially when clear evaluation criteria exist.
Example Applications:
  • Improving literary translations through multiple evaluation and refinement cycles.
  • Conducting multi-round research queries where additional iterations refine search results.
4.6 Taxonomy of Agentic RAG Systems
@@
4.6.1 Single-Agent Agentic RAG: Router
@@ 6 Tools and Frameworks for Agentic RAG
  • Semantic Kernel: Semantic Kernel [52, 53] is an open-source SDK by Microsoft that integrates large language models (LLMs) into applications. It supports agentic patterns, enabling the creation of autonomous AI agents for natural language understanding, task automation, and decision-making. It has been used in scenarios like ServiceNow’s P1 incident management to facilitate real-time collaboration, automate task execution, and retrieve contextual information seamlessly